# Instruction Tuning for Large Language Models: A Survey

### v5

@@ 8.1 HELM Evaluation
(1) Broad coverage. During the development, language models can be adapted to various NLP tasks (e.g., sequence labeling and question answering), thus, the evaluation of language models needs to be carried out in a wide range of scenarios. To involve all potential scenarios, HELM proposed a top-down taxonomy, which begins by compiling all existing tasks in a major NLP conference (ACL2022) into a task space and dividing each task into the form of scenarios (e.g., languages) and metrics (e.g., accuracy). Then when facing a specific task, the taxonomy would select one or more scenarios and metrics in the task space to cover it. By analyzing the structure of each task, HELM clarifies the evaluation content (task scenarios and metrics) and improves the scenario coverage of language models from 17.9% to 96.0%.
(2) Multi-metric measurement. In order to enable human to weigh language models from different perspectives, HELM proposes multi-metric measurement. HELM has covered 16 different scenarios and 7 metrics. To ensure the results of intensive multi-metric measurement, HELM measured 98 of 112 possible core scenarios (87.5%).
(3) Standardization. The increase in the scale and training complexity of language models has seriously hindered human’s understanding of the structure of each language model. To establish a unified understanding of existing language models, HELM benchmarks 30 well-known language models, covering such institutions as Google (UL2Tay et al. (2022)), OpenAI (GPT-3Brown et al. (2020b)), and EleutherAI (GPT-NeoXBlack et al. (2022)). Interestingly, HELM pointed out that LMs such as T5 Raffel et al. (2019) and Anthropic-LMv4-s3 Bai et al. (2022a) had not been directly compared in the initial work, while LLMs such as GPT-3 and YaLM were still different from their corresponding reports after multiple evaluations.
@@ 8.3 Smaller Instruction Dataset
IT requires a substantial amount of specialized instruction data for training. Zhou et al. (2023) hypothesized that the pre-trained LLM only has to learn the style or format to interact with users and proposed LIMA that achieves strong performance by fine-tuning an LLM on only 1,000 carefully selected training examples.
Specifically, LIMA first manually curates 1,000 demonstrations with high-quality prompts and responses. Then the 1,000 demonstrations are used to fine-tune the pre-trained 65B-parameter LLaMa Touvron et al. (2023b). By comparison, across more than 300 challenging tasks, LIMA outperfrms GPT-davinci003 Brown et al. (2020b), which was fine-tuned on 5,200 examples by human feedback tuning. Moreover, with only half amount of demonstrations, LIMA achieves equivalent results to GPT-4 OpenAI (2023), Claude Bai et al. (2022b), and Bard171717Bard, designed by Google, is an interface to generative AI platform, and the link is: https://ai.google/static/documents/google-about-bard.pdf. Above all, LIMA demonstrated that LLMs’ powerful knowledge and capabilities can be exposed to users with only a few carefully curated instructions to fine-tune.
8.4 Evaluating Instruction-tuning Datasets
@@
To address this issue, Wang et al. (2023c) performs dataset evaluation by fine-tuning the LLaMa model Touvron et al. (2023b) on various of open IT datasets and measure different fine-tuned models through both automatic and human evaluations. An additional model is trained on the combination of IT datasets. For the results, Wang et al. (2023c) showed that there is not a single best IT dataset across all tasks, while by manually combining datasets it can achieve the best overall performance. Besides, Wang et al. (2023c) pointed out that though IT can bring large benefits on LLMs at all sizes, smaller models and models with a high base quality benefit most from IT. For human evaluations, Wang et al. (2023c) a larger model is more likely to gain a higher acceptability score.
8.5 Do IT just learn Pattern Copying?
To address the lack of clarity about the specific knowledge that models acquire through instruction tuning, Kung and Peng (2023) delves into the analysis of how models make use of instructions during IT by comparing the tuning when provided with altered instructions versus the original instructions.
@@ 8.5 Do IT just learn Pattern Copying?
In summary, according to Kung and Peng (2023), the notable performance improvements observed in current IT models may be attributed to their ability to capture surface-level patterns, such as learning the output format and making guesses, rather than comprehending and learning the specific task.
@@ 9 Conclusion
This work surveys recent advances in the fast growing field of instruction tuning. We make a systematic review of the literature, including the general methodology of IT, the construction of IT datasets, the training of IT models, IT’s applications to different modalities, domains and application. We also review analysis on IT models to discover both their advantages and potential pitfalls. We hope this work will act as a stimulus to motivate further endeavors to address the deficiencies of current IT models.
@@ Appendix A Datasets
https://github.com/allenai/unifiedqa
@@
https://github.com/LAION-AI/Open-Instruction-Generalist
@@
https://github.com/hkunlp/unifiedskg
@@
https://github.com/allenai/natural-instructions-v1
@@
https://github.com/allenai/natural-instructions
@@
https://huggingface.co/datasets/bigscience/P3
@@
https://github.com/bigscience-workshop/xmtf
@@
https://github.com/google-research/FLAN
@@
https://github.com/BAAI-Zlab/COIG
@@
https://github.com/orhonovich/unnatural-instructions
@@
https://github.com/yizhongw/self-instruct
@@
https://github.com/XueFuzhao/InstructionWild
@@
https://github.com/nlpxucan/evol-instruct
@@
https://github.com/tatsu-lab/stanford_alpaca
@@
https://github.com/csitfun/LogiCoT
@@
https://huggingface.co/datasets/databricks/databricks-dolly-15k
@@
https://github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM
@@
https://huggingface.co/datasets/GAIR/lima
@@
https://huggingface.co/datasets/JosephusCheung/GuanacoDataset
@@
https://github.com/LAION-AI/Open-Assistant
@@
https://github.com/project-baize/baize-chatbot
@@
https://github.com/thunlp/UltraChat#data
@@
https://huggingface.co/datasets/Open-Orca/OpenOrca
@@
https://huggingface.co/datasets/RyokoAI/ShareGPT52K
@@
https://huggingface.co/datasets/allenai/WildChat
@@
https://github.com/ise-uiuc/magicoder?tab=readme-ov-file#-dataset
@@
https://huggingface.co/microsoft/phi-1
@@
https://huggingface.co/datasets/berkeley-nest/Nectar
@@
https://github.com/uclaml/SPIN?tab=readme-ov-file#Data

### v7

@@ 8.3 HELM Evaluation
(1) Broad coverage.
During the development, language models can be adapted to various NLP tasks (e.g., sequence labeling and question answering), thus, the evaluation of language models needs to be carried out in a wide range of scenarios. To involve all potential scenarios, HELM proposed a top-down taxonomy, which begins by compiling all existing tasks in a major NLP conference (ACL2022) into a task space and dividing each task into the form of scenarios (e.g., languages) and metrics (e.g., accuracy). Then when facing a specific task, the taxonomy would select one or more scenarios and metrics in the task space to cover it. By analyzing the structure of each task, HELM clarifies the evaluation content (task scenarios and metrics) and improves the scenario coverage of language models from 17.9% to 96.0%.
(2) Multi-metric measurement.
In order to enable human to weigh language models from different perspectives, HELM proposes multi-metric measurement. HELM has covered 16 different scenarios and 7 metrics. To ensure the results of intensive multi-metric measurement, HELM measured 98 of 112 possible core scenarios (87.5%).
(3) Standardization.
The increase in the scale and training complexity of language models has seriously hindered human’s understanding of the structure of each language model. To establish a unified understanding of existing language models, HELM benchmarks 30 well-known language models, covering such institutions as Google (UL2Tay et al. (2022)), OpenAI (GPT-3Brown et al. (2020b)), and EleutherAI (GPT-NeoXBlack et al. (2022)). Interestingly, HELM pointed out that LMs such as T5 Raffel et al. (2019) and Anthropic-LMv4-s3 Bai et al. (2022a) had not been directly compared in the initial work, while LLMs such as GPT-3 and YaLM were still different from their corresponding reports after multiple evaluations.
@@ 8.5 Smaller Instruction Dataset
IT requires a substantial amount of specialized instruction data for training. Zhou et al. (2023a) hypothesized that the pre-trained LLM only has to learn the style or format to interact with users and proposed LIMA that achieves strong performance by fine-tuning an LLM on only 1,000 carefully selected training examples.
Specifically, LIMA first manually curates 1,000 demonstrations with high-quality prompts and responses. Then the 1,000 demonstrations are used to fine-tune the pre-trained 65B-parameter LLaMa Touvron et al. (2023b). By comparison, across more than 300 challenging tasks, LIMA outperfrms GPT-davinci003 Brown et al. (2020b), which was fine-tuned on 5,200 examples by human feedback tuning. Moreover, with only half amount of demonstrations, LIMA achieves equivalent results to GPT-4 OpenAI (2023), Claude Bai et al. (2022b), and Bard191919Bard, designed by Google, is an interface to generative AI platform, and the link is: https://ai.google/static/documents/google-about-bard.pdf. Above all, LIMA demonstrated that LLMs’ powerful knowledge and capabilities can be exposed to users with only a few carefully curated instructions to fine-tune.
8.6 Evaluating Instruction Tuning Datasets
@@
To address this issue, Wang et al. (2023d) performs dataset evaluation by fine-tuning the LLaMa model Touvron et al. (2023b) on various of open IT datasets and measure different fine-tuned models through both automatic and human evaluations. An additional model is trained on the combination of IT datasets. For the results, Wang et al. (2023d) showed that there is not a single best IT dataset across all tasks, while by manually combining datasets it can achieve the best overall performance. Besides, Wang et al. (2023d) pointed out that though IT can bring large benefits on LLMs at all sizes, smaller models and models with a high base quality benefit most from IT. For human evaluations, Wang et al. (2023d) a larger model is more likely to gain a higher acceptability score.
8.7 Superficial Alignment
Despite the impressive improvements in the performance of instruction tuning, there lacks clarity about the specific knowledge that models acquire through instruction tuning, raising questions about: Does instruction tuning just learn Pattern Copying? or How exactly does the alignment tuning transform a base LLM?
To answer these questions, Kung and Peng (2023) delves into the analysis of how models make use of instructions during IT by comparing the tuning when provided with altered instructions versus the original instructions.
@@ 8.7 Superficial Alignment
Similar to the findings of Kung and Peng (2023), several subsequent studies Zhou et al. (2023a); Lin et al. (2023a) reached the same conclusion: the observed performance improvements in current IT models are often due to superficial alignment. This means the models excel at recognizing superficial alignment, such as mastering output formats and making educated guesses, rather than truly understanding and learning the underlying tasks.
@@ 9 Conclusion
This work surveys recent advances in the fast growing field of instruction tuning, which can also be referred to as supervised fine-tuning (SFT). We make a systematic review of the literature, including the general methodology of IT, the construction of IT datasets, the training of IT models, IT’s applications to different modalities, domains and application. We also review analysis on IT models to discover both their advantages and potential pitfalls. We hope this work will act as a stimulus to motivate further endeavors to address the deficiencies of current IT models.
@@ Appendix A Datasets
1 https://github.com/allenai/unifiedqa
@@
2 https://github.com/LAION-AI/Open-Instruction-Generalist
@@
3 https://github.com/hkunlp/unifiedskg
@@
4 https://github.com/allenai/natural-instructions-v1
@@
5 https://github.com/allenai/natural-instructions
@@
6 https://huggingface.co/datasets/bigscience/P3
@@
7 https://github.com/bigscience-workshop/xmtf
@@
8 https://github.com/google-research/FLAN
@@
9 https://github.com/BAAI-Zlab/COIG
@@
10 https://github.com/orhonovich/unnatural-instructions
@@
11 https://github.com/yizhongw/self-instruct
@@
12 https://github.com/XueFuzhao/InstructionWild
@@
13 https://github.com/nlpxucan/evol-instruct
@@
14 https://github.com/tatsu-lab/stanford_alpaca
@@
15 https://github.com/csitfun/LogiCoT
@@
16 https://huggingface.co/datasets/databricks/databricks-dolly-15k
@@
17 https://github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM
@@
18 https://huggingface.co/datasets/GAIR/lima
@@
19 https://huggingface.co/datasets/JosephusCheung/GuanacoDataset
@@
20 https://github.com/LAION-AI/Open-Assistant
@@
21 https://github.com/project-baize/baize-chatbot
@@
22 https://github.com/thunlp/UltraChat#data
@@
23 https://huggingface.co/datasets/Open-Orca/OpenOrca
@@
24 https://huggingface.co/datasets/RyokoAI/ShareGPT52K
@@
25 https://huggingface.co/datasets/allenai/WildChat
@@
26 https://github.com/ise-uiuc/magicoder?tab=readme-ov-file#-dataset
@@
27 https://huggingface.co/microsoft/phi-1
@@
28 https://huggingface.co/datasets/berkeley-nest/Nectar
@@
29 https://github.com/uclaml/SPIN?tab=readme-ov-file#Data